# Policy Optimization finds Nash Equilibrium in Regularized General-Sum LQ Games

### v1

@@ I Introduction
Recent years have witnessed a significant increase in the number of empirical works on reinforcement learning (RL) focusing on multi-agent competitive environments, in applications such as advertising [1], autonomous driving [2], multi-armed bandits [3], healthcare [4], and finance [5] among others. In many of these real-world problems [4, 5], guaranteed convergence and sample efficiency are critical due to the high cost of incorrect decisions. Hence, considerable theoretical attention has been directed towards specific instances of games, such as (1) purely cooperative games [6], (2) zero-sum games [7], (3) potential games [8, 9], or (4) (coarse) correlated equilibria in general-sum games [10, 11]. However, learning the Nash equilibrium (NE) for general-sum games is considered to be a significantly harder problem [7]. Recently, a burgeoning area of research has emerged [12], focusing on the development of provable Nash finding using RL techniques tailored for large population games through the mean-field game (MFG) paradigm [13, 14, 15], but they only provide approximate solutions for finite agent scenarios.
@@
    We characterize the NE of the ERGS game and find them to exclusively belong to the class of Linear Gaussian policies defined by a sequence of Coupled Riccati equations (Theorem III.1). Furthermore, we find that if the entropy regularization parameter $\tau$ is lower bounded by a model dependent scalar, then the uniqueness of the NE is guaranteed.
@@ III Nash Equilibrium Characterization
Further, we define the Nash value of the same agent $i$ as the cost incurred under a NE policy $J^{i*}_{t}(x)=J^{i}_{t}(x,\pi^{i*},\pi^{-i*})$. This helps us characterize the NE of the ERGS game (1)-(2) using the following theorem.
Theorem III.1.
Proof.
We start by proving a helping lemma.
Lemma III.2.
@@
Having proved the lemma we now proceed by proposing that the form of Nash value function (3) is quadratic, which we will prove by induction. Since we know that $J^{i*}_{T}(x)=x^{\top}Q^{i}_{T}x$, let
@@
Using Lemma III.2, we deduce that the NE policy of agent $i$ is given as
@@
satisfies the conditions (14) and (15), and hence, also the necessary conditions for the NE. Moreover, since the cost functional in (13) is strictly convex (due to Lemma III.2), we see that (16)-(17) is also sufficient for the NE.
@@
The proof of Theorem III.1 uses HJI equations [33] which reduce the problem to finding the most optimal probability distribution for each agent in each timestep given that the other agents follow NE policies. This simplification coupled with the fact that entropy regularized quadratic function is minimized by a Linear Guassian probability distribution (Lemma III.2), helps us arrive at the conclusion.
Theorem III.1 also states that although there may be multiple NE (any solution to (4) will specify a NE), all NE must have a linear Gaussian structure given by (6). Since all policies belong to the class of linear Gaussian distributions, we will henceforth restrict our discussion to policies of the form $S^{i}_{t}=(K^{i}_{t},\Sigma^{i}_{t})$ with $K^{i}_{t}\in\mathbb{R}^{p\times m}$ and $\Sigma^{i}_{t}\in\mathbb{R}^{p\times p}$ where $\Sigma^{i}_{t}$ is a symmetric positive definite matrix, such that $\pi^{i}_{t}(x)=u^{i}_{t}\sim\mathcal{N}(K^{i}_{t}x,\Sigma^{i}_{t})$. We also denote the joint policy at time $t\in\{0,\ldots,T-1\}$ as $S_{t}:=(S^{i}_{t})_{i\in[N]}$ and the joint policy for agent $i\in[N]$ as $S^{i}:=(S^{i}_{t})_{0\leq t\leq T-1}$. We further denote the set of all linear Gaussian policies for agent $i$ as $\mathcal{S}^{i}_{t}$, joint policies at time $t$ by $\mathcal{S}_{t}$ and the set of all joint policies for agent $i$ by $\mathcal{S}^{i}$. In keeping with game theoretic notation, superscript $-i$ denotes the set of all policies except for agent $i$; in particular, we have that $\mathcal{S}^{-i}=(\mathcal{S}^{j})_{j\in[N]\setminus\{i\}}$.
@@
Under this lower bound assumption on $\tau$, there exists a unique NE of the game (1)-(2) as proved below.
@@
The proof of Lemma III.4 relies on the invertibility of the $\Phi_{t}$ matrix at each timestep $t$, similar to the invertibility conditions in non-regularized LQ games [17], but relies on the entropy regularization parameter. Having proved the sufficient conditions for the existence and uniqueness of NE (6), next we move to PO methods for computation of the same.
@@ IV Policy Optimization (PO) & Non-Asymptotic Analysis
Using Theorem III.1, let us also define the best response map of agent $i$ at time $t$ to the set of policies $S_{t}$ as $\Psi^{i}_{t}(\cdot):\mathcal{S}_{t}\rightarrow\mathcal{S}^{i}_{t}$ such that
@@
$=\sum_{i\in[N]}\lVert K^{i}_{t}-K^{i\prime}_{t}\rVert_{F}+\lVert\Sigma^{i}_{t}-\Sigma^{i\prime}_{t}\rVert_{F}$
@@
Having formulated the auxiliary game, which is a LQ game, we can utilize [18, Theorem 4.5], which relies on the good convergence of the inner loop (lines 3-6 in Algorithm 1). This good convergence of the inner-loop is guaranteed under $\tau>2\gamma^{2}_{B}(\gamma^{*}_{P}+c\epsilon)(N-1)$ since [18, Theorem 4.5] inductively proves that if at time $t+1$, $\lVert P^{i}_{t+1}-P^{i*}_{t+1}\rVert=\mathcal{O}(\epsilon)$ and the inner loop is contractive (which is true under $\tau>2\gamma^{2}_{B}(\gamma^{*}_{P}+c\epsilon)(N-1)$), then $\lVert P^{i}_{t}-P^{i*}_{t}\rVert=\mathcal{O}(\epsilon)$. Since $\lVert P^{i}_{t}-P^{i*}_{t}\rVert=\mathcal{O}(\epsilon)$ and if $\epsilon>0$ is small enough, $\lVert\Sigma^{i}_{t}-\Sigma^{i*}_{t}\rVert=\mathcal{O}(\epsilon)$ where $\Sigma^{i}_{t}=\big{(}I+2(R^{i}_{t}+(B^{i}_{t})^{\top}P^{i}_{t+1}B^{i}_{t})/\tau\big{)}^{-1}$. The proof is thus complete. ∎
@@ IV-A $\delta$-augmented Entropy-Regularized Game
Now we investigate the case where the entropy regularization parameter is augmented with $\delta\in\mathbb{R}^{+}$, i.e., $\ddot{\tau}:=\tau+\delta$ is the entropy regularized parameter of the augmented game with dynamics (1) and cost function
@@
Furthermore, we assume that $\delta$ is chosen such that $\ddot{\tau}>2\gamma^{2}_{B}\ddot{\gamma}^{*}_{P}(N-1)$ where $\ddot{\gamma}^{*}_{P}:=\max_{\forall i,t}\lVert\ddot{P}^{i*}_{t}\rVert_{F}$ and $\ddot{P}^{i*}_{t}$ are solutions to the coupled Riccati equations of the augmented problem. We denote the unique NE of the augmented game by $(\ddot{S}^{i*}_{t})_{\forall i,t}=(\ddot{K}^{i*}_{t},\ddot{\Sigma}^{i*}_{t})_{\forall i,t}$ which is bound to exist due to the condition on $\ddot{\tau}$.
@@
$J^{i}(x,\ddot{S}^{i*},\ddot{S}^{-i*})\!\leq\!J^{i}(x,S^{i},\ddot{S}^{-i*})\!+\!\mathcal{O}(\delta),\forall i\in[N],S^{i}\in\mathcal{S}^{i}.$ (22)
@@
$\hskip 39.83368pt+(K^{i}_{s})^{\top}((\nicefrac{{\ddot{\tau}}}{{2}})I+R^{i}_{s})K^{i}_{s}\big{)}x_{s}+x^{\top}_{T}Q^{i}_{T}x_{T}\Big{|}x_{t}=x\Big{]},$
@@
which results in (22), which completes the proof. ∎

### v2

@@ I Introduction
Recent years have witnessed a significant increase in the number of empirical works on reinforcement learning (RL) focusing on multi-agent competitive environments, in applications such as autonomous driving [1], multi-armed bandits [2], healthcare [3], and finance [4] among others. In many of these real-world problems [3, 4], guaranteed convergence and sample efficiency are critical due to the high cost of incorrect decisions. Hence, considerable theoretical attention has been directed towards specific instances of games, such as (1) purely cooperative games [5], (2) zero-sum games [6], (3) potential games [7, 8], or (4) (coarse) correlated equilibria in general-sum games [9, 10]. However, learning the Nash equilibrium (NE) for general-sum games is considered to be a significantly harder problem [6]. Recently, a burgeoning area of research has emerged [11], focusing on the development of provable Nash finding using RL techniques tailored for large population games through the mean-field game (MFG) paradigm [12, 13, 14], but they only provide approximate solutions for finite agent scenarios.
@@
    We characterize the NE of the ERGS game and find them to exclusively belong to the class of Linear Gaussian policies defined by a sequence of Coupled Riccati equations (Theorem III.2). Furthermore, we find that if the entropy regularization parameter $\tau$ is lower bounded by a model dependent scalar, then the uniqueness of the NE is guaranteed.
@@ III Nash Equilibrium Characterization
Further, we define the Nash value of the same agent $i$ as the cost incurred under a NE policy $J^{i*}_{t}(x)=J^{i}_{t}(x,\pi^{i*},\pi^{-i*})$. Before characterizing the NE of the ERGS game we provide a helper lemma which might be of independent interest.
Lemma III.1.
@@
This lemma helps us characterize the NE of the ERGS game (1)-(2) presented in the following theorem.
Theorem III.2.
Proof.
We start by proposing that the form of Nash value function (3) is quadratic, which we will prove by induction. Since we know that $J^{i*}_{T}(x)=x^{\top}Q^{i}_{T}x$, let
@@
Using Lemma III.1, we deduce that the NE policy of agent $i$ is given as
@@
satisfies the conditions (14) and (15), and hence, also the necessary conditions for the NE. Moreover, since the cost functional in (13) is strictly convex (due to Lemma III.1), we see that (16)-(17) is also sufficient for the NE.
@@
The proof of Theorem III.2 uses HJI equations [32] which reduce the problem to finding the most optimal probability distribution for each agent (in each timestep) given that the other agents follow NE policies. This simplification coupled with the fact that entropy regularized quadratic function is minimized by a Linear Guassian probability distribution (Lemma III.1), helps us arrive at the conclusion.
Theorem III.2 also states that although there may be multiple NE (any solution to (9) will specify a NE), all NE must have a linear Gaussian structure given by (11). Since all policies belong to the class of linear Gaussian distributions, we will henceforth restrict our discussion to policies of the form $S^{i}_{t}=(K^{i}_{t},\Sigma^{i}_{t})$ with $K^{i}_{t}\in\mathbb{R}^{p\times m}$ and $\Sigma^{i}_{t}\in\mathbb{R}^{p\times p}$ where $\Sigma^{i}_{t}$ is a symmetric positive definite matrix, such that $\pi^{i}_{t}(x)=u^{i}_{t}\sim\mathcal{N}(K^{i}_{t}x,\Sigma^{i}_{t})$. We also denote the joint policy at time $t\in\{0,\ldots,T-1\}$ as $S_{t}:=(S^{i}_{t})_{i\in[N]}$ and the joint policy for agent $i\in[N]$ as $S^{i}:=(S^{i}_{t})_{0\leq t\leq T-1}$. We further denote the set of all linear Gaussian policies for agent $i$ as $\mathcal{S}^{i}_{t}$, joint policies at time $t$ by $\mathcal{S}_{t}$ and the set of all joint policies for agent $i$ by $\mathcal{S}^{i}$. In keeping with game theoretic notation, superscript $-i$ denotes the set of all policies except for agent $i$; in particular, we have that $\mathcal{S}^{-i}=(\mathcal{S}^{j})_{j\in[N]\setminus\{i\}}$.
@@
This recursive assumption is similar in structure to the condition on attenuation parameter in robust control problems [34, 35]. Under this lower bound assumption on $\tau$, there exists a unique NE of the game (1)-(2) as proved below.
@@
The proof of Lemma III.4 relies on the invertibility of the $\Phi_{t}$ matrix at each timestep $t$, similar to the invertibility conditions in non-regularized LQ games [16], but relies on the entropy regularization parameter. Having established sufficient conditions for the existence and uniqueness of NE (11), next we move to PO methods for computation of the same.
@@ IV Policy Optimization (PO) & Non-Asymptotic Analysis
Using Theorem III.2, let us also define the best response map of agent $i$ at time $t$ to the set of policies $S_{t}$ as $\Psi^{i}_{t}(\cdot):\mathcal{S}_{t}\rightarrow\mathcal{S}^{i}_{t}$ such that
@@
Having formulated the auxiliary game, which is a LQ game, we can utilize [17, Theorem 4.5], which relies on the good convergence of the inner loop (lines 3-6 in Algorithm 1). This good convergence of the inner-loop is guaranteed under $\tau>2\gamma^{2}_{B}(\gamma^{*}_{P}+c\epsilon)(N-1)$ since [17, Theorem 4.5] inductively proves that if at time $t+1$, $\lVert P^{i}_{t+1}-P^{i*}_{t+1}\rVert=\mathcal{O}(\epsilon)$ and the inner loop is contractive (which is true under $\tau>2\gamma^{2}_{B}(\gamma^{*}_{P}+c\epsilon)(N-1)$), then $\lVert P^{i}_{t}-P^{i*}_{t}\rVert=\mathcal{O}(\epsilon)$. Since $\lVert P^{i}_{t}-P^{i*}_{t}\rVert=\mathcal{O}(\epsilon)$ and if $\epsilon>0$ is small enough, $\lVert\Sigma^{i}_{t}-\Sigma^{i*}_{t}\rVert=\mathcal{O}(\epsilon)$ where $\Sigma^{i}_{t}=\big{(}I+2(R^{i}_{t}+(B^{i}_{t})^{\top}P^{i}_{t+1}B^{i}_{t})/\tau\big{)}^{-1}$. The proof is thus complete. ∎
@@ IV-A $\delta$-augmented Entropy-Regularized Game
Now we investigate the case where the entropy regularization parameter $\tau$ does not satisfy Assumption III.3 leading to possible non-uniqueness and indeterminacy of NE. Let us define an augmented game with an augmented entropy regularized parameter $\ddot{\tau}:=\tau+\delta$, dynamics (1) and cost function
@@
where $\delta\in\mathbb{R}^{+}$ is chosen such that $\ddot{\tau}>2\gamma^{2}_{B}\ddot{\gamma}^{*}_{P}(N-1)$ where $\ddot{\gamma}^{*}_{P}:=\max_{\forall i,t}\lVert\ddot{P}^{i*}_{t}\rVert_{F}+c\epsilon$ ($c$ and $\epsilon$ similar to Theorem IV.2) and $\ddot{P}^{i*}_{t}$ are solutions to the coupled Riccati equations of the augmented problem. Using results in the previous sections the NE of the augmented game exists, is unique and can be found using Algorithm 1. Furthermore, as shown below the NE of the augmented game will be $\mathcal{O}(\delta)$-NE of the original game. We denote the unique NE of the augmented game by $(\ddot{S}^{i*}_{t})_{\forall i,t}=(\ddot{K}^{i*}_{t},\ddot{\Sigma}^{i*}_{t})_{\forall i,t}$.
@@
$J^{i}(x,\ddot{S}^{i*},\ddot{S}^{-i*})\!\leq\!J^{i}(x,S^{i},\ddot{S}^{-i*})\!+\!\mathcal{O}(\delta),\forall i\in[N],S^{i}\in\mathcal{S}^{i}.$
@@
$\hskip 34.14322pt+(K^{i}_{s})^{\top}((\nicefrac{{\ddot{\tau}}}{{2}})I+R^{i}_{s})K^{i}_{s}\big{)}x_{s}+x^{\top}_{T}Q^{i}_{T}x_{T}\Big{|}x_{t}=x\Big{]},$
@@
which completes the proof. ∎